# L4Q: Parameter Efficient Quantization-Aware Fine-Tuning on Large Language Models

### v4

@@ 5 Conclusion
In this work, we introduce L4Q, a parameter-efficient quantization-aware fine-tuning method for large language models. L4Q enables element-wise adaptation of model weights for downstream tasks while simultaneously optimizing quantization parameters. This concurrent optimization ensures that the adaptation parameters effectively account for quantization errors. We demonstrate the efficiency of L4Q, which significantly reduces training resource requirements compared to traditional QAT. Moreover, since the L4Q layer is designed to produce fully quantized low-bit model weights, it maintains inference efficiency, unlike QLoRA, LoftQ, or QAT-LoRA, which result in mixed-precision models. The effectiveness of L4Q as a QAT framework is further supported by experimental results across various task evaluations. L4Q consistently achieves superior quality maintenance on language tasks, demonstrating its enhanced adaptability compared to QAT-LoRA and PTQ-based PEFT methods.
Ethics Statements
This paper presents work aimed at advancing the field of machine learning, particularly in the context of language models. While our research contributes to the development and exploitation of LLMs, we acknowledge that there are potential societal consequences associated with this work. However, we do not delve into the potential harms or biases inherent in these models.
Reproducibility Statement
This paper introduces a novel algorithm for quantization layer design and the training process. To ensure reproducibility, the source code is provided as an anonymous, downloadable link in the supplementary materials, along with detailed explanations of the environmental setup and execution process. The theoretical results are thoroughly explained with clear assumptions, and complete proofs are included in the appendix. Additionally, detailed descriptions of the data processing steps for all datasets used in the experiments are available in the supplementary materials. These resources are intended to fully enable the reproduction of our results.
@@ Appendix A Details on Quantization Scale Learning Procedure
A.1 Quantization Scale update on QAT
From the conditions and notations in Equation 3, Equations  4 and  5 are derived as follows. First, the derivative of $s$ is presented as follows.
$\frac{\partial W_{q}}{\partial s}=\frac{\partial}{\partial s}(\tilde{w}\times s+b)=s\frac{\partial}{\partial s}(\tilde{w})+\tilde{w}=s\frac{\partial}{\partial s}(R\cdot clamp(w))+\tilde{w}$ (13)
By applying the STE, the rounding function $R$ is considered as an identity function. Therefore, the rounding function, combined with a clamp function $\tilde{R}:=R\cdot clamp$, and its derivative is induced as follows. Note that $w=\frac{W-b}{s}$.
$\tilde{R}(w)=\begin{cases}Qn,&if\enspace w<Qn\\ w,&if\enspace Q_{n}\leq w\leq Q_{p}\\ Qp,&if\enspace w>Q_{p}\end{cases}\quad\frac{\partial}{\partial w}\tilde{R}(w)=\begin{cases}1,&if\enspace Q_{n}\leq w\leq Q_{p}\\ 0,&otherwise\end{cases}$ (14)
By applying the chain rule, the derivation of term $R\cdot clamp(w)=\tilde{R}((W-b)/s)$ is expressed as below.
$\frac{\partial}{\partial s}(\tilde{R}(w))=\frac{\partial\tilde{R}}{\partial w}\frac{\partial w}{\partial s}=\frac{\partial\tilde{R}}{\partial w}\frac{\partial}{\partial s}(\frac{W-b}{s})=\frac{\partial\tilde{R}}{\partial w}(-\frac{W-b}{s^{2}})$ (15)
@@ A.1 Quantization Scale update on QAT
$\frac{\partial W_{q}}{\partial s}=s\frac{\partial\tilde{R}}{\partial w}(-\frac{W-b}{s^{2}})+\tilde{w}=\frac{\partial\tilde{R}}{\partial w}(-\frac{W-b}{s})+\tilde{w}=\begin{cases}Qn,&if\enspace w<Qn\\ -w+\tilde{w},&if\enspace Q_{n}\leq w\leq Q_{p}\\ Qp,&if\enspace w>Q_{p}\end{cases}$ (16)
@@
$\frac{\partial W_{q}}{\partial b}$ $=\frac{\partial}{\partial b}(\tilde{w}\times s+b)=s\frac{\partial}{\partial b}(\tilde{R}(w))+1=s\frac{\partial\tilde{R}}{\partial w}(\frac{\partial}{\partial b}(\frac{W-b}{s}))+1$ (17)
$=\frac{\partial\tilde{R}}{\partial w}(-1)+1=\begin{cases}0,&if\enspace Q_{n}\leq w\leq Q_{p}\\ 1,&otherwise\end{cases}$ (18)
@@
A.2 Quantization scale and LoRA Parameter update on L4Q
@@
$w=\frac{W_{0}+\alpha BA-b}{s},\enspace\tilde{w}=\tilde{R}(w)\quad s.t.\quad\tilde{R}=R\cdot clamp$ (20)
@@
$\frac{\partial w}{\partial A}=\frac{\alpha B^{\top}}{s},\quad\frac{\partial w}{\partial B}=\frac{\alpha A^{\top}}{s},\quad\frac{\partial W}{\partial w}=\frac{\partial}{\partial w}(\tilde{R}(w)s+b)=s\frac{\partial\tilde{R}}{\partial w}$ (23)
Therefore, by substitution of Equation 23 and applying STE on $\frac{\partial\tilde{R}}{\partial w}$ from Equation 14 on Equation  22, the equation is simplified by the crossed-out products between the terms. As a result, the partial derivatives presented in Equation 10 can be derived as follows.
$\frac{\partial W_{q}}{\partial A}=\frac{\partial w}{\partial A}\frac{\partial W_{q}}{\partial w}=(\frac{\alpha B^{\top}}{s})(\frac{s\partial\tilde{R}}{\partial w})=\begin{cases}\alpha B^{\top},&if\enspace Q_{n}\leq w\leq Q_{p}\\ 0,&otherwise\end{cases}$ (24)
$\frac{\partial W_{q}}{\partial B}$ $=\frac{\partial W}{\partial w}\frac{\partial w}{\partial B}=(s\frac{\partial\tilde{R}}{\partial w})(\frac{\alpha A^{\top}}{s})=\begin{cases}\alpha A^{\top},&if\enspace Q_{n}\leq w\leq Q_{p}\\ 0,&otherwise\end{cases}$ (25)
Finally, substitution of Equation 24 and  25 to Equation 21 derives the Equation 26 and  27.
@@ Appendix B Quantization initialization
We report the detailed model accuracy evaluation results of L4Q fine-tuning across different initialization methods, along with the quantization error and clipping error for each method, measured both at the initialization point and the end of the training, in Table  4. The LLaMA-2 7B model was trained for a total of 12,800 iterations with batch size 128, using the same hyperparameters as those in the main evaluation.
@@
Appendix C Effect of LoRA Warm-up on L4Q training
We investigated the effect of LoRA warm-up on L4Q training and found that a small number of LoRA warm-up steps is often beneficial for quickly restoring model accuracy. Using LLaMA-1 and LLaMA-2 7B models, we trained the models for a total of 12,800 iterations with 128 batches. The remaining training conditions, such as the quantization target layers and the rank of LoRA parameters, were kept consistent with the main experimental setup. The Commonsense QA evaluation results are presented in Table 5.
Table 5: Commonsense QA benchmark result with the variations of LoRA warm-up on LLaMA-1 and LLaMA-2 7B models. The numbers represent the accuracy(%) of each task.
Model #Bits Warmup Hella. PIQA ARC-c ARC-e Winog. BoolQ OBQA Avg.
LLaMA-1 7B 4 0 58.0 78.9 45.1 76.6 70.1 76.5 37.0 63.2
10 58.0 78.5 44.9 76.4 69.7 76.5 36.4 62.9
20 57.9 78.8 45.2 76.4 70.1 76.1 36.2 62.9
40 57.5 78.6 45.6 76.0 69.8 75.8 35.4 62.7
80 57.2 78.6 45.1 75.3 69.6 75.1 34.8 62.2
3 0 55.8 77.2 41.9 73.6 68.1 76.1 32.8 60.8
10 55.7 77.2 42.0 74.0 68.2 75.9 33.0 60.8
20 55.0 77.5 41.5 74.0 68.0 74.9 32.2 60.4
40 54.3 77.3 41.1 73.3 66.5 73.7 31.8 59.7
80 52.6 76.8 39.5 71.7 65.4 71.0 31.8 58.4
LLaMA-2 7B 4 0 57.5 78.1 46.9 75.8 69.8 76.0 35.0 62.7
10 57.1 78.2 46.0 76.0 69.5 75.8 35.8 62.6
20 57.2 78.5 46.1 76.5 69.9 77.7 34.2 62.8
40 57.1 78.2 45.7 76.5 69.5 77.3 34.8 62.7
80 56.5 78.4 46.2 76.4 70.7 74.5 34.0 62.4
3 0 55.6 77.9 42.1 74.3 69.5 72.0 34.0 60.8
10 55.7 78.6 44.0 74.2 68.4 72.5 34.8 61.2
20 55.5 77.8 43.0 74.6 69.0 71.7 34.2 60.8
40 55.4 77.8 43.4 73.6 68.9 71.7 35.0 60.8
80 53.9 76.8 42.1 73.7 67.3 71.8 34.4 60.0
Models with fewer than 20 LoRA warm-up steps performed best, as initializing LoRA parameters with minimal disturbance from quantization errors facilitated better convergence. Extending the LoRA warm-up beyond the quantization-aware fine-tuning phase, however, resulted in performance degradation, likely due to an insufficient number of training steps to fully compensate for quantization errors. Considering the variability in impact across different models and configurations, we generally applied 10 LoRA warm-up steps in our experiments.
Appendix D Ablative Study on LoRA Rank
We investigated the effect of LoRA rank size on L4Q training and found that a rank size of around 4 is sufficient for effective training. Using the LLaMA-1 7B model, we conducted training over 12,800 iterations with 128 batches. The remaining training conditions are consistent with the main experiments. The evaluation results for Commonsense QA and MMLU are presented in Table 6 and Table 7, respectively.
Table 6: Commonsense QA benchmark result on LLaMA-1 7B model. The numbers represent accuracy (%) for each task.
@@ Appendix D Ablative Study on LoRA Rank
LLaMA-1 7B 1 56.9 79.1 43.7 76.0 69.9 76.6 35.4 62.5
2 58.6 78.6 45.4 76.8 69.9 74.1 36.4 62.8
4 57.8 79.1 45.3 76.0 69.5 76.1 34.8 62.7
8 58.5 79.1 45.8 76.5 70.2 75.8 37.0 62.7
16 58.3 78.8 44.5 75.8 69.8 78.3 35.2 63.0
32 58.6 79.0 45.2 76.3 70.2 76.7 37.6 63.4
64 58.4 79.0 45.5 76.5 70.0 78.7 37.8 63.2
128 58.6 79.1 45.9 76.7 70.6 75.8 36.4 63.3
Table 7: MMLU benchmark result on LLaMA-1 7B model. The numbers represent accuracy (%) for each category.
@@
LLaMA-1 7B 1 31.8 29.2 32.9 37.0 32.7 32.3 30.7 35.2 39.9 34.4
2 30.7 28.3 32.7 37.5 32.2 33.0 29.3 34.3 37.2 33.5
4 32.4 32.1 36.7 39.4 34.9 32.9 31.4 38.7 39.5 35.4
8 33.5 31.9 37.7 39.4 35.5 34.2 30.7 38.4 39.8 35.7
16 30.0 31.6 35.1 38.1 33.5 32.2 31.2 35.3 38.1 34.8
32 32.1 30.8 33.2 36.5 33.1 32.5 31.7 37.2 39.5 35.0
64 33.4 31.2 36.1 38.8 34.8 33.6 31.1 35.9 39.5 35.0
128 31.8 29.7 34.5 37.6 33.3 33.3 31.1 35.7 38.3 34.5
Increasing the rank beyond 32 or 4 does not lead to further performance improvements, which aligns with the observations in the original LoRA paper (Hu et al., 2022). Therefore, we generally applied a rank size of 4, considering that higher rank sizes introduce memory and computational overhead during training.
Appendix E Experimental Settings
The baselines and L4Q are trained with AdamW optimizer Loshchilov & Hutter (2019) with a weight decay of 0.01. For the learning rate scheduler, a cosine decay scheduler with a linear warm-up through 10% of the total training steps. Learning rates are presented in Table 8.
@@ Methods
LLaMA-2 7B $2\times 10^{-5}$ $2\times 10^{-4}$ $2\times 10^{-4}$
LLaMA-2 13B $2\times 10^{-5}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Batch size is set to 128. 50K iterations for the baselines that utilize PTQ applied schemes, such as QLoRA* and QA-LoRA, and 25K iterations for QAT involving schemes, such as QAT, QAT-LoRA, and L4Q are utilized to match the training latency. The training sequence length of training is set to above the maximum sequence length of the dataset, which is 2048, except for a 33B model with L4Q that is set to 128.
Appendix F Training Time
We report the total training time of L4Q and the quantization-aware PEFT baselines, QLoRA* and QA-LoRA, in Table 9. L4Q demonstrates similar time performance to the baselines, highlighting its scalability and time efficiency for larger model sizes, comparable to that of the baseline methods.

### v5

@@ 5 Conclusion
In this work, we introduce L4Q, a parameter-efficient quantization-aware fine-tuning method for large language models. L4Q enables element-wise adaptation of model weights for downstream tasks while simultaneously optimizing quantization parameters. This concurrent optimization ensures that the adaptation parameters effectively account for quantization errors. We demonstrate the efficiency of L4Q, which significantly reduces training resource requirements compared to traditional QAT. Moreover, since the L4Q layer is designed to produce fully quantized low-bit model weights, it maintains inference efficiency, unlike QLoRA, LoftQ, or QAT-LoRA, which results in mixed precision models. The effectiveness of L4Q as a QAT framework is further supported by experimental results in various task evaluations. L4Q consistently achieves superior quality maintenance in language tasks, demonstrating its enhanced adaptability compared to the QAT-LoRA and PTQ-based PEFT methods.
Limitations
Our work focuses on efficient weight quantization methods for large language models (LLMs), but there are additional avenues that could further enhance inference efficiency and effectiveness, particularly when combined with L4Q. Activation quantization offers a promising way to further reduce memory and computation costs when combined with weight quantization, potentially optimizing LLM performance. Similarly, KV cache compression through pruning or quantization could minimize memory overhead and latency, especially for long-context applications. Finally, refinement of LoRA initialization schemes for quantized models may enhance fine-tuning efficiency and improve performance in low-resource settings.
Although we have not investigated these vertical methods within the context of L4Q, we believe that these directions offer promising opportunities for future research. Integrating these approaches with L4Q could further improve LLM inference efficiency and effectiveness, which we leave to future work.
Ethical Considerations
This paper presents work aimed at advancing machine learning, particularly in the context of language models. While our research contributes to the development and application of LLMs, we recognize the potential societal implications associated with this work. Based on the claims of the referenced sources — including datasets, code, and models — we believe that the artifacts used and the models produced do not violate personal identification rights or contain offensive content. Consequently, we do not further examine or discuss potential harms or biases inherent in these models.
All datasets, code, and models cited in this paper are publicly accessible and processed solely for research purposes. Our use of these artifacts is consistent with their intended purpose.
@@ Appendix A Details on Quantization Scale Learning Procedure
A.1 Quantization parameter update on QAT
From the conditions and notations in Equation 3, Equations  5 and  6 are derived as follows. First, the derivative of $s$ is presented as follows.
$\frac{\partial W_{q}}{\partial s}=\frac{\partial}{\partial s}(\tilde{w}\times s+b)=s\frac{\partial}{\partial s}(\tilde{w})+\tilde{w}=s\frac{\partial}{\partial s}(\text{round}\cdot\text{clamp}(w))+\tilde{w}$ (16)
By applying the STE, the rounding function is considered as an identity function. Therefore, the rounding function, combined with a clamp function $\tilde{\text{r}}:=\text{round}\cdot\text{clamp}$, and its derivative is induced as follows. Note that $w=\frac{W-b}{s}$.
$\tilde{\text{r}}(w)=\begin{cases}Qn,&if\enspace w<Qn\\ w,&if\enspace Q_{N}\leq w\leq Q_{P}\\ Qp,&if\enspace w>Q_{P}\end{cases}\quad\frac{\partial}{\partial w}\tilde{\text{r}}(w)=\begin{cases}1,&if\enspace Q_{N}\leq w\leq Q_{P}\\ 0,&otherwise\end{cases}$ (17)
By applying the chain rule, the derivation of term $\tilde{\text{r}}(w)=\tilde{\text{r}}((W-b)/s)$ is expressed as below.
$\frac{\partial}{\partial s}(\tilde{\text{r}}(w))=\frac{\partial\tilde{\text{r}}}{\partial w}\frac{\partial w}{\partial s}=\frac{\partial\tilde{\text{r}}}{\partial w}\frac{\partial}{\partial s}(\frac{W-b}{s})=\frac{\partial\tilde{\text{r}}}{\partial w}(-\frac{W-b}{s^{2}})$ (18)
@@ A.1 Quantization parameter update on QAT
$\frac{\partial W_{q}}{\partial s}=s\frac{\partial\tilde{\text{r}}}{\partial w}(-\frac{W-b}{s^{2}})+\tilde{w}=\frac{\partial\tilde{\text{r}}}{\partial w}(-\frac{W-b}{s})+\tilde{w}=\begin{cases}Qn,&if\enspace w<Qn\\ -w+\tilde{w},&if\enspace Q_{N}\leq w\leq Q_{P}\\ Qp,&if\enspace w>Q_{P}\end{cases}$ (19)
@@
$\frac{\partial W_{q}}{\partial b}$ $=\frac{\partial}{\partial b}(\tilde{w}\times s+b)=s\frac{\partial}{\partial b}(\tilde{\text{r}}(w))+1=s\frac{\partial\tilde{\text{r}}}{\partial w}(\frac{\partial}{\partial b}(\frac{W-b}{s}))+1$ (20)
$=\frac{\partial\tilde{\text{r}}}{\partial w}(-1)+1=\begin{cases}0,&if\enspace Q_{N}\leq w\leq Q_{P}\\ 1,&otherwise\end{cases}$ (21)
@@
A.2 Quantization parameter and LoRA Parameter update on L4Q
@@
$w=\frac{W_{0}+\alpha BA-b}{s},\enspace\tilde{w}=\tilde{\text{r}}(w)\quad s.t.\quad\tilde{\text{r}}=\text{round}\cdot\text{clamp}$ (23)
@@
$\frac{\partial w}{\partial A}=\frac{\alpha B^{\top}}{s},\quad\frac{\partial w}{\partial B}=\frac{\alpha A^{\top}}{s},\quad\frac{\partial W}{\partial w}=\frac{\partial}{\partial w}(\tilde{\text{r}}(w)s+b)=s\frac{\partial\tilde{\text{r}}}{\partial w}$ (26)
Therefore, by substitution of Equation 26 and applying STE on $\frac{\partial\tilde{\text{r}}}{\partial w}$ from Equation 17 on Equation  25, the equation is simplified by the crossed-out products between the terms. As a result, the partial derivatives presented in Equation 12 can be derived as follows.
$\frac{\partial W_{q}}{\partial A}=\frac{\partial w}{\partial A}\frac{\partial W_{q}}{\partial w}=(\frac{\alpha B^{\top}}{s})(\frac{s\partial\tilde{\text{r}}}{\partial w})=\begin{cases}\alpha B^{\top},&if\enspace Q_{N}\leq w\leq Q_{P}\\ 0,&otherwise\end{cases}$ (27)
$\frac{\partial W_{q}}{\partial B}$ $=\frac{\partial W}{\partial w}\frac{\partial w}{\partial B}=(s\frac{\partial\tilde{\text{r}}}{\partial w})(\frac{\alpha A^{\top}}{s})=\begin{cases}\alpha A^{\top},&if\enspace Q_{N}\leq w\leq Q_{P}\\ 0,&otherwise\end{cases}$ (28)
Finally, substitution of Equation 27 and  28 to Equation 24 derives the Equation 29 and  30.
@@ Appendix B Quantization initialization
We report the detailed model accuracy evaluation results of L4Q fine-tuning across different initialization methods, along with the quantization error and clipping error for each method, measured both at the initialization point and the end of the training.The LLaMA-2 7B model was trained for 12,800 iterations with a batch size of 128, using the same hyperparameters as in the main evaluation.
As shown in Table 4, while the overall quantization error remains relatively consistent across initialization methods and fine-tuned states, the clipping error exhibits significant variation. The clipping error reflects the number of values clipped during quantization, and different initialization methods lead to varying degrees of clipping throughout training. Notably, L4Q achieves the lowest clipping error and the highest model accuracy, demonstrating the effectiveness of its initialization strategy.
@@
Appendix C Ablative Study on L4Q Hyperparameters
C.1 LoRA Rank Size
We investigated the effect of LoRA rank size on L4Q training. Using the LLaMA-2 7B model, we conducted training over 12,800 iterations with 128 batches. The remaining training conditions are consistent with the main experiments. The evaluation results for CSQA and MMLU are presented in Table 5 and Table 6, respectively.
Table 5: Commonsense QA benchmark result on LLaMA-2 7B model. The numbers represent accuracy (%) for each task.
@@ C.1 LoRA Rank Size
LLaMA-2 7B 1 57.6 78.4 45.5 77.0 68.2 77.4 34.0 62.6
2 57.4 78.5 45.1 76.2 69.3 77.6 34.0 62.6
4 57.5 78.2 46.1 77.1 68.7 78.1 35.4 63.0
8 56.9 78.5 46.3 78.1 69.3 77.8 34.8 63.1
16 57.2 78.1 46.3 77.2 68.7 78.9 33.4 62.8
32 57.8 78.4 46.2 77.1 68.7 78.2 35.8 63.2
64 57.4 78.6 46.1 77.1 69.5 78.5 34.8 63.1
128 57.5 78.1 46.0 77.0 68.8 78.4 35.4 63.0
Table 6: MMLU benchmark result on LLaMA-2 7B model. The numbers represent accuracy (%) for each category.
@@
LLaMA-2 7B 1 37.4 33.3 43.4 44.0 39.4 40.8 36.8 48.2 49.2 43.5
2 38.1 31.6 41.6 42.2 38.4 42.2 35.2 48.6 49.0 43.7
4 36.3 33.4 42.7 43.2 38.7 42.5 36.6 50.2 51.2 44.9
8 39.5 34.6 45.0 45.0 41.0 42.7 36.7 50.3 51.7 45.0
16 38.7 35.8 45.7 45.8 41.3 42.0 36.6 49.4 49.7 44.3
32 39.4 35.0 46.1 45.6 41.3 42.4 37.1 49.8 49.0 44.4
64 39.6 35.0 45.8 47.2 41.7 43.6 37.4 50.9 50.9 45.6
128 38.8 35.4 44.8 44.7 40.7 42.8 36.6 50.3 50.6 44.9
Increasing the rank beyond 4 does not lead to significant performance improvements, which aligns with the observations in the original LoRA paper (Hu et al., 2022). Therefore, we generally applied a rank size of 4, considering that higher rank sizes introduce memory and computational overhead during training.
C.2 Quantization Group Size
We investigated the effect of quantization group size on L4Q training. Using the LLaMA-2 7B model, we conducted experiments with group size 32 to 128 with the same training conditions in the main experiments. The evaluation results for Commonsense QA and MMLU are presented in Table 7 and Table 8, respectively.
Table 7: Commonsense QA benchmark result on LLaMA-2 7B model. The numbers represent accuracy (%) for each task.
Model Group Size Hellaswag PIQA ARC-c ARC-e Winogrande BoolQ OBQA Average
LLaMA-2 7B 128 57.2 78.8 47.1 76.9 70.2 80.4 34.8 63.6
64 57.5 77.5 46.7 78.3 70.2 80.7 34.8 63.7
32 57.6 77.6 47.6 78.2 70.1 80.6 35.4 63.9
Table 8: MMLU benchmark result on LLaMA-2 7B model. The numbers represent accuracy (%) for each category.
0-shot 5-shot
Model Group Size Hums. STEM Social Others Avg. Hums. STEM Social Others Avg.
LLaMA-2 7B 128 38.7 33.8 45.6 46.4 40.9 42.9 37.7 50.5 51.9 45.5
64 38.2 35.6 47.2 46.5 41.5 43.8 37.0 52.2 52.7 46.3
32 39.5 35.9 47.8 46.4 42.1 44.4 38.2 52.3 52.5 46.7
Having a fine-grained quantization group size leads to performance improvements, which aligns with the observations in the conventional group-wise quantization works (Frantar et al., 2023). We applied a quantization with the group size of 128 considering that smaller quantization group sizes introduce memory and computational overhead during inference and training.
Appendix D Experimental Settings
The baselines and L4Q are trained with AdamW optimizer Loshchilov and Hutter (2019) with a weight decay of 0.01. For the learning rate scheduler, a cosine decay scheduler with a linear warm-up through 10% of the total training steps. Learning rates are presented in Table 9.
@@ Methods
LLaMA-2 7B $2\times 10^{-5}$ $2\times 10^{-5}$ $2\times 10^{-4}$ $2\times 10^{-4}$
LLaMA-2 13B $2\times 10^{-5}$ $2\times 10^{-5}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Mistral-v0.1 7B $1\times 10^{-5}$ $5\times 10^{-6}$ $5\times 10^{-6}$
The batch size is set to 128. For baselines that utilize PTQ-based schemes, such as QLoRA* and QA-LoRA, training is conducted for 50K iterations. For QAT-based methods, such as QAT, QAT-LoRA, and L4Q, training is conducted for 25K iterations. This reduction in training length for QAT-based methods is due to their faster convergence, as illustrated in Figure 5 with an example of LLaMA-2 7B.
Using the same training hyperparameters, including a learning rate of $2\times 10^{-5}$, the joint training of quantization parameters and LoRA weight parameters enables L4Q to converge more quickly. This allows for halving the training length, which also helps mitigate overfitting.
Additionally, the training sequence length is set to match or exceed the maximum sequence length of the dataset, which is 2048. The only exception is the 33B model with L4Q, where the training sequence length is set to 128.
Figure 5: Train loss curve of L4Q and QLoRA. With a same training condition, L4Q converges faster than QLoRA.
Appendix E Training Efficiency
E.1 Memory Usage
We report the memory usage of QAT (LSQ), quantization-aware PEFT baselines (LoftQ, QLoRA, QA-LoRA, QAT-LoRA), and L4Q in Table 10. Note that QLoRA reduces memory usage further by employing paged optimizers, a technique that can also be applied to other fine-tuning methods, including L4Q. We plan to explore this vertical implementation in future work. As quantization-aware PEFT methods utilize pre-quantized model weights, the memory usage of L4Q and quantization-aware PEFT methods differs by the amount of the reduced memory usage of model weights during inference.
OpenLLaMA LLaMA
Methods 3B 7B 13B 33B
LoRA 15.1 25.1 43.8 71.9
LoftQ 5.2 7.9 19.6 31.9
QLoRA 5.2 7.9 19.6 31.9
QA-LoRA 7.8 14.8 27.8 67.2
LSQ 44.2 79.5 OOM OOM
QAT-LoRA 22.6 41.9 70.6 OOM
L4Q 15.3 25.4 44.3 73.2
Table 10: Memory cost (GB) for fine-tuning LLMs on a single NVIDIA A100 GPU. (OOM: Out of Memory)
E.2 Training Time
We report the total training time of L4Q, QAT (LSQ), and the quantization-aware PEFT baselines(QLoRA* and QA-LoRA) in Table 11. While L4Q has a longer training time per step compared to the baselines due to gradient recomputation during the backpropagation stage, the reduced number of training steps enables L4Q to achieve similar overall training time performance. The training time for QAT on 13B and 33B models was measured using 2 and 4 NVIDIA A100 GPUs on a single node.